Miltmar Corporation will pay a year-end dividend of $5, and dividends thereafter are expected to grow at the constant rate of 6% per year. The risk-free rate is 5%, and the expected return on the market portfolio is 10%. The stock has a beta of 0.76.

a.

Calculate the market capitalization rate. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  Market capitalization rate %
b.

What is the intrinsic value of the stock? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  Intrinsic value $   
Answer #1

a). Market Capitalization Rate = k = risk-free rate + [beta * (expected market return - risk-free rate)]

= 5% + [0.76 * (10% - 5%)] = 5% + 3.80% = 8.80%

b). Intrinsic Value = V0 = Next Dividend / (k - growth rate)

= $5 / (0.0880 - 0.06) = $5 / 0.028 = $178.57